Suppose household consumption is C=100+0.90Y, where Y=income, and investment is 150. Find the equilibrium output and consumption.
1
Expert's answer
2015-01-21T02:58:26-0500
Answer on Question #50382, Economics - Macroeconomics

Assignment

Suppose household consumption is C=100+0.90Y, whereY=income, and investment is 150. Find the equilibrium output and consumption.

Solution

We can write an equation
Y = Equilibrium output = C + I
Y = 100+0.90Y + 150
0.1Y = 250
Y = 2500
Now we can find consumption
C= Consumption = 100+0.90Y
C=100+0.90*2500
C = 2350